OpenStudy (anonymous):

What is the range of y = 3sinx + 3?

OpenStudy (anonymous):

[-3, 3]

OpenStudy (anonymous):

What is the domain of y = tan(x) + 2?

OpenStudy (anonymous):

x is not equal to pi/2 + n(pi) where n is any integer

OpenStudy (across):

The range of \(y=\sin(x)\) is \([-1,1]\). The range of \(y=3\sin(x)\) is \([-3,3]\). The range of \(y=3\sin(x)+3\) is \([0,6]\).\[\]

OpenStudy (anonymous):

domain is all real numbers except odd-integer multiples of pi/2.